OpenStudy (anonymous):

How Many three digit numbers can be written using no 0's and at least one 7 answer is 217 but why?

OpenStudy (cwrw238):

possibilities are xx7 x7x 7xx 777 77x 7x7 x77 where x = 1-6, 8 or 9

OpenStudy (zarkon):

(# ways to have 3digit number with no zeros)-(#ways to have 3 digit number with no zeros or sevens)

OpenStudy (cwrw238):

ahhh - yes

OpenStudy (zarkon):

cwrw238's way will also work

OpenStudy (anonymous):

@AlexBabich is repetition allowed

OpenStudy (anonymous):

no every number can only be used once

OpenStudy (zarkon):

if it is 'at least one 7' then I'd say repetition is allowed

OpenStudy (zarkon):

otherwise it would say you have to use a 7

OpenStudy (cwrw238):

yes

OpenStudy (zarkon):

also...with repetition you get the answer given in the OP

OpenStudy (anonymous):

what do you mean by repitition?

OpenStudy (zarkon):

777 is a possibility...557 is a possibility...ect

OpenStudy (anonymous):

ahh yea of course, I thought it was meant repeating the same number

OpenStudy (cwrw238):

for xx7, number of xx's will be 8P2 + 8 right?

OpenStudy (anonymous):

no just 8*8

OpenStudy (zarkon):

8P2 + 8 =8*8

OpenStudy (anonymous):

what does 8p2 mean, its not 8 power 2?

OpenStudy (zarkon):

8P2=8*7

OpenStudy (cwrw238):

number of permutations of 2 in 8 = 8*7

OpenStudy (zarkon):

\[_nP_r=\frac{n!}{(n-r)!}\]

OpenStudy (cwrw238):

so i guess that would be 64 ways for x7x and 7xx total * 64 = 192

OpenStudy (cwrw238):

and the others are 1 + 8 + 8 + 8 = 25

OpenStudy (anonymous):

yea i added up all the ways you gave me and got 217, but i didnt use permutations, I havent even learned that, i just did 8*8 for the ones with 2 spaces added them up and did...what you just typed is what i did

OpenStudy (zarkon):

9*9*9-8*8*8

OpenStudy (cwrw238):

- that IS an easier way!!

OpenStudy (anonymous):

@Zarkon yes i did the way you said and got the same result

OpenStudy (anonymous):

Thank you both so much! :)

OpenStudy (cwrw238):

yw